Andrew saved his coins for several weeks. Before the carnival,he noticed he had twice as many quarters as dimes,and 8 more dimes than nickels. How much money did Andrew have?

Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

Before the carnival,he noticed he had twice as many quarters as dimes,and 8 more dimes than nickels. How much money did Andrew have?
--------
q = 2d
d = n + 8
------
Total = 10d + 25q + 5n cents
Total = 10d + 25(2d) + 5(d-8) cents
Total = 10d + 50d + 5d - 40
Total = 65d - 40
Total = 650-40 = 610 cents = $6.10
-------------
Cheers,
Stan H.
-----------

Answer by MathTherapy(10810)


Andrew saved his coins for several weeks. Before the carnival,he noticed he had twice as many quarters as dimes,and 8 more dimes than nickels. How much money did Andrew have?
Are you serious? How can one tell?